OpenStudy (anonymous):

Solve log3 27 - x = 5 Question 8 options: 1) x = -216 2) x = -2 3) x = 8 4) No Solution

OpenStudy (dumbcow):

x=-2

OpenStudy (anonymous):

log3 27-x=5 3log3 3-x=5 3-x=5 x=-2